Given that the Young modulus of steel is 210 GPa, the extension of a steel wire undergoing a stress of 300x106 Pa is equal to:

A) 0.14%

B) 1.4%

C) 14%

D) 140%


1
Expert's answer
2022-03-27T10:56:39-0400
"\sigma=\varepsilon E"

"\varepsilon =\frac{\sigma}{E}=\frac{300*10^6}{210*10^9}=0.0014=0.14\%"

Answer: A) 0.14%